# Dynamical disequilibrium in dwarf galaxies: rethinking gas dynamics, rotation curves, and dark matter inference

### v1

@@ Abstract
pubyear: 2025pagerange: Dynamical disequilibrium in dwarf galaxies: rethinking gas dynamics, rotation curves, and dark matter inferenceB
@@ 1 Introduction
Dwarf galaxies – broadly defined as systems with stellar masses $M_{\star}<10^{9}\,\mathrm{M}_{\odot}$ – fall in a crucial regime for testing both galaxy formation physics and the nature of dark matter. As the most dark-matter-dominated galaxies known, they (potentially) offer clean laboratories for studying gravitational dynamics, and thus dark matter, in low baryon-fraction environments (see e.g. Sales2023). At the same time, their shallow potential wells render them highly susceptible to feedback processes, introducing complex baryonic effects that complicate both simulations and observations (e.g., baryon-induced core creation, BICC, Pontzen2012; Llambay2019).
@@
Recent advances have shifted the focus from inner slopes to more global discrepancies. Cosmological parameters are now tightly constrained on large scales (e.g. Planck2016), and dark matter halo scaling relations – such as the dependence of halo structural properties on cosmology – are well established (e.g. Correa2015c; Ludlow2016; DiemerJoyce2019). This enables robust predictions of halo circular velocity profiles once a single parameter (e.g., the maximum circular velocity, $V_{\rm max}$) is known.
@@
Following this logic, Oman2015 reformulated the cusp–core problem in terms of an inner mass deficit: a shortfall in the circular velocity at a fixed inner radius (e.g. 2 kpc, see also Santos2020, for a later, more flexible definition) relative to that expected from the dark matter halo alone. Some galaxies exhibit rotation velocities far below this baseline, even when sharing the same $V_{\rm max}$ as systems without such deficits. The phenomenon spans a wide range of galaxy masses, from dwarfs to large discs, and varies substantially at fixed galaxy properties (e.g., the baryonic surface mass density $\Sigma_{\rm bar}$, or the baryonic mass contribution in the inner regions $\eta_{\mathrm{bar}}=(V_{\rm bar,fid}/V_{\rm fid})^{2}$).
@@
In extreme cases, the inferred deficits exceed the total baryonic mass, challenging the hypothesis that they arise from BICC (Oman2015). While such processes – driven by bursty feedback and gravitational coupling (Navarro_Eke_Frenk1996; Mashchenko2008; Governato2012; Pontzen2012) – can flatten central density profiles, simulations cannot reproduce the most extreme cores inferred from observations and indicate that these baryonic processes are limited to a narrow range of halo masses (DiCintio2014; Chan2015). Moreover, not all state-of-the-art simulations predict cores: cosmological hydrodynamical simulations like EAGLE (Schaye2014; Schaller2015, see also Llambay2019) and Illustris (Vogelsberger2014; Chua2019) reproduce key galaxy statistics without featuring them.
@@
In addition to a solution based on modifications due to baryonic feedback, alternative explanations fall into two main categories, as outlined by Oman2015: (i) dark matter is not cold and collisionless; or (ii) mass distributions are incorrectly inferred from observations.
@@
Departures from the central assumption of the cold dark matter framework – that dark matter behaves as a cold, collisionless fluid – can alter the internal kinematics of dwarf galaxies. Self-interacting dark matter (SIDM) models have been proposed (Spergel2000), in which scattering interactions among DM particles are introduced. For interaction cross-sections per unit mass of order $\sigma/m=1~\mathrm{cm^{2}~g^{-1}}$, such interactions can thermalize the central regions of DM haloes, producing the constant-density cores observed in dwarfs while retaining the large-scale successes of CDM (see e.g. Rocha2013; Elbert2015; Ren2019). SIDM has also been invoked to explain the broad diversity of dwarf galaxy rotation curves, however, additional effects due to galaxy formation physics complicate this picture: introducing baryon-DM coupling and fine-tuning seems inevitable (Creasey2017; Kaplinghat2020; Burger2022; Correa2025).
@@
Nevertheless, the possibility that the observed cores are merely artifacts of systematic uncertainties in observation or modelling has not been conclusively ruled out. Dark matter haloes have long been predicted to be triaxial in shape, as indicated by some of the earliest CDM simulations (Davis1985; Frenk88). Such triaxiality gives rise to an aspherical gravitational potential, with the strongest deviations from spherical symmetry occurring in the central regions (Hayashi2007). A disc galaxy embedded within such a halo – typically aligned with the plane defined by the halo’s intermediate and major axes (Hayashi2006) – thus experiences a non-axisymmetric potential. This induces non-circular motions in the disc, as gas orbits become elongated, producing primarily bisymmetric ($m=2$ harmonic) variations in azimuthal velocity as a function of projection angle (e.g. Marasco2018; Oman2019). Even mild asphericity in the potential can generate appreciable non-circular motions (e.g. Hayashi2006). Although harmonic decompositions of observed dwarf galaxy velocity fields generally indicate small non-circular motion amplitudes (Trachternach2008), these are likely underestimated in standard model fitting procedures (Chemin2020). In many cases, velocity variations arising from non-circular motions are absorbed by degenerate model parameters (e.g. Roper2023), potentially leading to substantial errors in rotation curve measurements – that is, significant discrepancies between the inferred and true rotational velocities at a given radius.
Despite the promising efforts to resolve the diversity of rotation curve shapes tension in studies targeting the effects of non-circular motions (Marasco2018; Oman2019; Roper2023) and/or pressure gradients (e.g. Valenzuela2007; Chemin2016; Read2016; Pina2025), they have yet to provide conclusive evidence that the tension has been fully resolved (Santos2020; see also Sales2023 for a recent review). This realisation has motivated a closer examination of the hierarchy of assumptions required to reduce the complex dynamics of gaseous discs to a one-dimensional rotation curve – i.e. an azimuthally averaged rotational velocity profile, $V_{\rm rot}(r)$ – as a tracer of the underlying gravitational potential.
@@
Recent simulation studies have shown that gas in dwarf galaxies often violates these equilibrium assumptions, owing to a range of perturbative processes. Bursty feedback, accretion flows, misaligned angular momentum, and environmental interactions can generate warps, asymmetries, and large-scale non-circular motions (e.g. ElBadry2018; DowningOman2023; Sands2024). These effects challenge the validity of conventional 3D or 2D velocity-field modelling pipelines, even when sophisticated corrections to the rotation curve are applied, as such methods implicitly assume steady-state dynamics.
Following this motivation, we present a novel, fully local, Eulerian analysis of midplane gas dynamics in a large sample of dwarf galaxies drawn from the high-resolution COLIBRE cosmological hydrodynamical simulations (Schaye2025; Chaikin2025). COLIBRE directly and self-consistently models the multiphase interstellar medium (ISM), including dust grains coupled to the chemical network. It reproduces key galaxy statistics at $z=0$, such as the observed stellar mass function, sizes, star formation rates, Hi and H2 masses, gas and stellar metallicities, and dust abundances (Schaye2025). These successes, and in particular the self-consistent treatment of the cold and dense phases of the ISM and the super-sampling of dark matter particles (hence enhanced resolution of the gravitational potential, mitigating spurious baryon–DM energy transfer Ludlow2019a; Ludlow2021; Ludlow2023), make COLIBRE particularly well suited to identifying where steady-state assumptions break down, and to tracing the physical origins of such departures from equilibrium.
@@ 2.1 COLIBRE simulations
COLIBRE uses the sphenix hydrodynamics scheme (Borrow2022), a smoothed particle hydrodynamics (SPH) formulation optimized for galaxy formation. This method employs a density–energy approach with artificial viscosity and conduction terms, using a quartic spline kernel with smoothing length $\eta=1.2348$, corresponding to approximately 65 weighted neighbours.
@@
Key subgrid physics modules in COLIBRE enhance the realism of the interstellar and circumgalactic media, allowing gas to cool to temperatures as low as $\sim$10 K. Radiative cooling, heating, and species abundances (atoms, ions, and molecules) are computed using hybrid-chimes (Richings2014a; Richings2014b; Ploeckinger2025), which evolves the non-equilibrium abundances of electrons and nine hydrogen and helium species on the fly, and treats an additional 147 metal species using quasi-equilibrium chemistry. Dust physics is modeled self-consistently for three grain species and two size bins, coupled to the chemistry and cooling network (Trayford2025). Star formation follows a Schmidt law based on gravitational instability, without imposing an effective ISM equation of state (Nobels2024).
@@
Stellar mass loss includes metal injection from asymptotic giant branch (AGB) stars, core-collapse supernovae (CCSNe), and type Ia supernovae (SNIa) with updated yields and a calibrated SNIa delay-time distribution (Correa at al., in preparation). Pre-supernova feedback from massive stars – including stellar winds, radiation pressure, and Hii region photoionization – is explicitly modelled (Llambay2025). CCSN and SNIa feedback follow a stochastic thermal scheme (DallaVecchia2012), supplemented by a lower-energy kinetic mode that enhances turbulence in the case of CCSN (Chaikin2023). The simulations analysed here employ the fiducial purely thermal active galactic nucleus (AGN) feedback model (BoothSchaye2009), updated to improve feedback sampling for low-mass black holes. We have explicitly checked that similar runs using the hybrid AGN model of Husko2025b – which tracks black hole spin and combines thermally driven winds with kinetic jet feedback – yield results consistent with the thermal AGN model in all parts of our analysis.
@@
COLIBRE employs four times more dark matter than baryonic particles, improving sampling of the gravitational potential and suppressing spurious CDM–baryon energy transfer (Ludlow2019a; Ludlow2021; Ludlow2023). The absence of an imposed ISM equation of state allows gas structure to develop on sub-kiloparsec scales, yielding a higher effective resolution than in previous simulations with comparable baryonic particle masses. Together, these features enable a fully multiphase ISM and improved modelling of the gravitational potential, representing a major resolution improvement over earlier galaxy formation simulations (see Schaye2025; Chaikin2025 for full model and calibration details, respectively).
@@
Haloes and galaxies are identified through a three-step process: (i) a friends-of-friends (FoF) algorithm applied to dark matter particles with a linking length of 0.2 times the mean interparticle spacing; (ii) the hydrodynamics-optimised hbt-herons subhalo finder (ForouharMoreno2025), a descendant of hbt+ (Han2018), which identifies bound structures and constructs merger trees; and (iii) the soap tool (McGibbon2025), which computes galaxy and halo properties within fixed and overdensity-based apertures, adopting $M_{200\mathrm{c}}$ and $R_{200\mathrm{c}}$ as the halo mass and radius definitions [^2] The subscript $200\mathrm{c}$ indicates a quantity evaluated within a spherical aperture with a mean enclosed density of $200$ times the critical density for closure, $\rho_{\mathrm{c}}=3H^{2}/8\pi G$.. Throughout the analysis, we refer to individual galaxies via their hbt-herons track ID, which uniquely links each $z=0$ halo to its main progenitor across simulation snapshots and thus provides a stable and physically meaningful identifier for each system.
@@ 2.2.1 Disc extent
The radial extent of each gas disc is defined using the cumulative Hi mass profile [^3] While not unique to COLIBRE, this approach is greatly facilitated by the self-consistent treatment of Hi fractions directly within the simulation outputs, avoiding the need for post-processing schemes commonly required in other simulations (see, e.g., the discussion in sec. 3.3 of Oman2019)., $M_{\rm H\textsc{i}}(<r)$. We identify the radius $R_{\rm ext,H\textsc{i}}$ at which this profile flattens, i.e., where the logarithmic slope satisfies
@@
with $\lambda=0.01~\mathrm{dex}~\mathrm{kpc}^{-1}$ chosen empirically to ensure robustness across the sample. For a rotationally supported disc this is usually similar to the radius enclosing 90 per cent of the Hi mass in a galaxy ($R_{90,\rm H\textsc{i}}$, see e.g. DowningOman2023; Sands2024); but unlike an enclosed mass threshold, it is a more reliable tracer of the outer edge of a galaxy’s gas in disturbed or irregular morphologies.
@@ 2.2.2 Warped disc frame
The coordinate origin is fixed at the potential minimum as identified by hbt-herons. The velocity centroid $\boldsymbol{v}_{\rm cen}$ is computed as the mass-weighted average velocity of dark matter particles within a sphere of radius $r_{\rm vel\,cen}=2\,r_{\rm conv}$, where $r_{\rm conv}$ is the convergence radius following Power2003. This choice ensures the frame is centered on a physically meaningful, dynamically stable origin and velocity frame, minimizing artificial harmonic modes (e.g., spurious $m=1$ modes) in the potential and velocity fields caused by miscentering.

### v2

@@ Abstract
pubyear: 2025pagerange: Dynamical disequilibrium in dwarf galaxies: rethinking gas dynamics, rotation curves, and dark matter inferenceC
@@ 1 Introduction
Dwarf galaxies – broadly defined as systems with stellar masses $M_{\star}<10^{9}\,\mathrm{M}_{\odot}$ – fall in a crucial regime for testing both galaxy formation physics and the nature of dark matter. As the most dark-matter-dominated galaxies known, they (potentially) offer clean laboratories for studying gravitational dynamics, and thus dark matter, in low baryon-fraction environments (see e.g. Sales et al., 2022). At the same time, their shallow potential wells render them highly susceptible to feedback processes, introducing complex baryonic effects that complicate both simulations and observations (e.g., baryon-induced core creation, BICC, Pontzen and Governato, 2012; Benítez-Llambay et al., 2019).
@@
Recent advances have shifted the focus from inner slopes to more global discrepancies. Cosmological parameters are now tightly constrained on large scales (e.g. Planck Collaboration et al., 2016), and dark matter halo scaling relations – such as the dependence of halo structural properties on cosmology – are well established (e.g. Correa et al., 2015; Ludlow et al., 2016; Diemer and Joyce, 2019). This enables robust predictions of halo circular velocity profiles once a single parameter (e.g., the maximum circular velocity, $V_{\rm max}$) is known.
@@
Following this logic, Oman et al. (2015) reformulated the cusp–core problem in terms of an inner mass deficit: a shortfall in the circular velocity at a fixed inner radius (e.g. 2 kpc, see also Santos-Santos et al., 2020, for a later, more flexible definition) relative to that expected from the dark matter halo alone. Some galaxies exhibit rotation velocities far below this baseline, even when sharing the same $V_{\rm max}$ as systems without such deficits. The phenomenon spans a wide range of galaxy masses, from dwarfs to large discs, and varies substantially at fixed galaxy properties (e.g., the baryonic surface mass density $\Sigma_{\rm bar}$, or the baryonic mass contribution in the inner regions $\eta_{\mathrm{bar}}=(V_{\rm bar,fid}/V_{\rm fid})^{2}$).
@@
In extreme cases, the inferred deficits exceed the total baryonic mass, challenging the hypothesis that they arise from BICC (Oman et al., 2015). While such processes – driven by bursty feedback and gravitational coupling (Navarro et al., 1996; Mashchenko et al., 2008; Governato et al., 2012; Pontzen and Governato, 2012) – can flatten central density profiles, simulations cannot reproduce the most extreme cores inferred from observations and indicate that these baryonic processes are limited to a narrow range of halo masses (Di Cintio et al., 2014; Chan et al., 2015). Moreover, not all state-of-the-art simulations predict cores: cosmological hydrodynamical simulations like EAGLE (Schaye et al., 2015; Schaller et al., 2015, see also Benítez-Llambay et al., 2019) and Illustris (Vogelsberger et al., 2014; Chua et al., 2019) reproduce key galaxy statistics without featuring them.
@@
In addition to a solution based on modifications due to baryonic feedback, alternative explanations fall into two main categories, as outlined by Oman et al. (2015): (i) dark matter is not cold and collisionless; or (ii) mass distributions are incorrectly inferred from observations.
@@
Departures from the central assumption of the cold dark matter framework – that dark matter behaves as a cold, collisionless fluid – can alter the internal kinematics of dwarf galaxies. Self-interacting dark matter (SIDM) models have been proposed (Spergel and Steinhardt, 2000), in which scattering interactions among DM particles are introduced. For interaction cross-sections per unit mass of order $\sigma/m=1~\mathrm{cm^{2}~g^{-1}}$, such interactions can thermalize the central regions of DM haloes, producing the constant-density cores observed in dwarfs while retaining the large-scale successes of CDM (see e.g. Rocha et al., 2013; Elbert et al., 2015; Ren et al., 2019). SIDM has also been invoked to explain the broad diversity of dwarf galaxy rotation curves, however, additional effects due to galaxy formation physics complicate this picture: introducing baryon-DM coupling and fine-tuning seems inevitable (Creasey et al., 2017; Kaplinghat et al., 2020; Burger et al., 2022; Correa et al., 2025).
@@
Nevertheless, the possibility that the observed cores are merely artifacts of systematic uncertainties in observation or modelling has not been conclusively ruled out. Dark matter haloes have long been predicted to be triaxial in shape, as indicated by some of the earliest CDM simulations (Davis et al., 1985; Frenk et al., 1988). Such triaxiality gives rise to an aspherical gravitational potential, with the strongest deviations from spherical symmetry occurring in the central regions (Hayashi et al., 2007). A disc galaxy embedded within such a halo – typically aligned with the plane defined by the halo’s intermediate and major axes (Hayashi and Navarro, 2006) – thus experiences a non-axisymmetric potential. This induces non-circular motions in the disc, as gas orbits become elongated, producing primarily bisymmetric ($m=2$ harmonic) variations in azimuthal velocity as a function of projection angle (e.g. Marasco et al., 2018; Oman et al., 2019). Even mild asphericity in the potential can generate appreciable non-circular motions (e.g. Hayashi and Navarro, 2006). Although harmonic decompositions of observed dwarf galaxy velocity fields generally indicate small non-circular motion amplitudes (Trachternach et al., 2008), these are likely underestimated in standard model fitting procedures (Chemin et al., 2020). In many cases, velocity variations arising from non-circular motions are absorbed by degenerate model parameters (e.g. Roper et al., 2023), potentially leading to substantial errors in rotation curve measurements – that is, significant discrepancies between the inferred and true rotational velocities at a given radius.
Despite the promising efforts to resolve the diversity of rotation curve shapes tension in studies targeting the effects of non-circular motions (Marasco et al., 2018; Oman et al., 2019; Roper et al., 2023) and/or pressure gradients (e.g. Valenzuela et al., 2007; Chemin et al., 2016; Read et al., 2016; Mancera Piña et al., 2025), they have yet to provide conclusive evidence that the tension has been fully resolved (Santos-Santos et al. 2020; see also Sales et al. 2022 for a recent review). This realisation has motivated a closer examination of the hierarchy of assumptions required to reduce the complex dynamics of gaseous discs to a one-dimensional rotation curve – i.e. an azimuthally averaged rotational velocity profile, $V_{\rm rot}(r)$ – as a tracer of the underlying gravitational potential.
@@
Recent simulation studies have shown that gas in dwarf galaxies often violates these equilibrium assumptions, owing to a range of perturbative processes. Bursty feedback, accretion flows, misaligned angular momentum, and environmental interactions can generate warps, asymmetries, and large-scale non-circular motions (e.g. El-Badry et al., 2018; Downing and Oman, 2023; Jahn et al., 2023; Rey et al., 2024; Sands et al., 2024). These effects challenge the validity of conventional 3D or 2D velocity-field modelling pipelines, even when sophisticated corrections to the rotation curve are applied, as such methods implicitly assume steady-state dynamics.
Following this motivation, we present a novel, fully local, Eulerian analysis of midplane gas dynamics in a large sample of dwarf galaxies drawn from the high-resolution COLIBRE cosmological hydrodynamical simulations (Schaye et al., 2025; Chaikin et al., 2025). COLIBRE directly and self-consistently models the multiphase interstellar medium (ISM), including dust grains coupled to the chemical network. It reproduces key galaxy statistics at $z=0$, such as the observed stellar mass function, sizes, star formation rates, Hi and H2 masses, gas and stellar metallicities, and dust abundances (Schaye et al., 2025). These successes, and in particular the self-consistent treatment of the cold and dense phases of the ISM and the super-sampling of dark matter particles (hence enhanced resolution of the gravitational potential, mitigating spurious baryon–DM energy transfer Ludlow et al., 2019, 2021, 2023), make COLIBRE particularly well suited to identifying where steady-state assumptions break down, and to tracing the physical origins of such departures from equilibrium.
@@ 2.1 COLIBRE simulations
COLIBRE uses the sphenix hydrodynamics scheme (Borrow et al., 2022), a smoothed particle hydrodynamics (SPH) formulation optimized for galaxy formation. This method employs a density–energy approach with artificial viscosity and conduction terms, using a quartic spline kernel with smoothing length $\eta=1.2348$, corresponding to approximately 65 weighted neighbours.
@@
Key subgrid physics modules in COLIBRE enhance the realism of the interstellar and circumgalactic media, allowing gas to cool to temperatures as low as $\sim$10 K. Radiative cooling, heating, and species abundances (atoms, ions, and molecules) are computed using hybrid-chimes (Richings et al., 2014a, b; Ploeckinger et al., 2025), which evolves the non-equilibrium abundances of electrons and nine hydrogen and helium species on the fly, and treats an additional 147 metal species using quasi-equilibrium chemistry. Dust physics is modeled self-consistently for three grain species and two size bins, coupled to the chemistry and cooling network (Trayford et al., 2025). Star formation follows a Schmidt law based on gravitational instability, without imposing an effective ISM equation of state (Nobels et al., 2024).
@@
Stellar mass loss includes metal injection from asymptotic giant branch (AGB) stars, core-collapse supernovae (CCSNe), and type Ia supernovae (SNIa) with updated yields and a calibrated SNIa delay-time distribution (Correa at al., in preparation). Pre-supernova feedback from massive stars – including stellar winds, radiation pressure, and Hii region photoionization – is explicitly modelled (Benítez-Llambay et al., 2025). CCSN and SNIa feedback follow a stochastic thermal scheme (Dalla Vecchia and Schaye, 2012), supplemented by a lower-energy kinetic mode that enhances turbulence in the case of CCSN (Chaikin et al., 2023). The simulations analysed here employ the fiducial purely thermal active galactic nucleus (AGN) feedback model (Booth and Schaye, 2009), updated to improve feedback sampling for low-mass black holes. We have explicitly checked that similar runs using the hybrid AGN model of Huško et al. (2025) – which tracks black hole spin and combines thermally driven winds with kinetic jet feedback – yield results consistent with the thermal AGN model in all parts of our analysis.
@@
COLIBRE employs four times more dark matter than baryonic particles, improving sampling of the gravitational potential and suppressing spurious CDM–baryon energy transfer (Ludlow et al., 2019, 2021, 2023). The absence of an imposed ISM equation of state allows gas structure to develop on sub-kiloparsec scales, yielding a higher effective resolution than in previous simulations with comparable baryonic particle masses. Together, these features enable a fully multiphase ISM and improved modelling of the gravitational potential, representing a major resolution improvement over earlier galaxy formation simulations (see Schaye et al. 2025; Chaikin et al. 2025 for full model and calibration details, respectively).
@@
Haloes and galaxies are identified through a three-step process: (i) a friends-of-friends (FoF) algorithm applied to dark matter particles with a linking length of 0.2 times the mean interparticle spacing; (ii) the hydrodynamics-optimised hbt-herons subhalo finder (Forouhar Moreno et al., 2025), a descendant of hbt+ (Han et al., 2018), which identifies bound structures and constructs merger trees; and (iii) the soap tool (McGibbon et al., 2025), which computes galaxy and halo properties within fixed and overdensity-based apertures, adopting $M_{200\mathrm{c}}$ and $R_{200\mathrm{c}}$ as the halo mass and radius definitions [^2] The subscript $200\mathrm{c}$ indicates a quantity evaluated within a spherical aperture with a mean enclosed density of $200$ times the critical density for closure, $\rho_{\mathrm{c}}=3H^{2}/8\pi G$.. Throughout the analysis, we refer to individual galaxies via their hbt-herons track ID, which uniquely links each $z=0$ halo to its main progenitor across simulation snapshots and thus provides a stable and physically meaningful identifier for each system.
@@ 2.2.1 Disc extent
The radial extent of each gas disc is defined using the cumulative Hi mass profile [^3] While not unique to COLIBRE, this approach is greatly facilitated by the self-consistent treatment of Hi fractions directly within the simulation outputs, avoiding the need for post-processing schemes commonly required in other simulations (see, e.g., the discussion in sec. 3.3 of Oman et al., 2019)., $M_{\rm H\textsc{i}}(<r)$. We identify the radius $R_{\rm ext,H\textsc{i}}$ at which this profile flattens, i.e., where the logarithmic slope satisfies
@@
with $\lambda=0.01~\mathrm{dex}~\mathrm{kpc}^{-1}$ chosen empirically to ensure robustness across the sample. For a rotationally supported disc this is usually similar to the radius enclosing 90 per cent of the Hi mass in a galaxy ($R_{90,\rm H\textsc{i}}$, see e.g. Downing and Oman, 2023; Sands et al., 2024); but unlike an enclosed mass threshold, it is a more reliable tracer of the outer edge of a galaxy’s gas in disturbed or irregular morphologies.
@@ 2.2.2 Warped disc frame
The coordinate origin is fixed at the potential minimum as identified by hbt-herons. The velocity centroid $\boldsymbol{v}_{\rm cen}$ is computed as the mass-weighted average velocity of dark matter particles within a sphere of radius $r_{\rm vel\,cen}=2\,r_{\rm conv}$, where $r_{\rm conv}$ is the convergence radius following Power et al. (2003). This choice ensures the frame is centered on a physically meaningful, dynamically stable origin and velocity frame, minimizing artificial harmonic modes (e.g., spurious $m=1$ modes) in the potential and velocity fields caused by miscentering.